# Physics-informed Kolmogorov–Arnold Network with Chebyshev Polynomials for Fluid Mechanics

### v2

@@ II.3 Physics-informed KAN Network
For each physical equation, we design a varying number of hidden layers from shallow to deep to compare and select the optimal number of hidden layers. The Adam optimizer is used for the parameter adjustments. Specific parameters are presented in Table 1, where ILR is the initial learning rate, and $Data_{train}$ and $Data_{test}$ are the number of samples that have been used for training and testing, respectively.
@@
Allen–Cahn equation 225 1000 [2] + [5] $\times$ (1$\sim$4) + [1] Adam $10^{-3}$ $10^{4}$ 0.1
Burgers equation 225 1000 [2] + [5] $\times$ (1$\sim$4) + [1] Adam $10^{-3}$ $10^{4}$ 0.1
Helmholtz equation 225 1000 [2] + [5] $\times$ (1$\sim$4) + [1] Adam $10^{-3}$ $10^{4}$ 0.1
Kovasznay flow 225 1500 [2] + [5] $\times$ (1$\sim$4) + [3] Adam $10^{-3}$ $10^{4}$ 0.1
cylinder wake flow 2000 5000 [3] + [5] $\times$ (1$\sim$4) + [3] Adam $10^{-3}$ $10^{4}$ 0.1
cavity flow 625 1200 [2] + [20/30/50] $\times$ (1$\sim$4) + [3] Adam $10^{-3}$ $2\times 10^{5}$ 0.1
@@
The two-dimensional input flow conditions are randomly sampled from a database of ground truth at a resolution of 15$\times$15 to form the training dataset, for more complex 2D cases like cavity flow, 25$\times$25 ground truth data is utilized as the training dataset. While such low-resolution data theoretically cannot fully describe the characteristics of the flow field, it more closely aligns with the sparsely discrete data obtained from real experiments through sensors. This presents a significant challenge to the inferential capabilities of the neural network, it forces the neural network to develop robust feature extraction capabilities under information-deficient conditions, better preparing it for practical deployment scenarios. For three-dimensional flow conditions, the amount of real data used for network training is appropriately increased.
@@ III Physical Models and Training Process of ChebPIKAN Models
For each equation, we explored six different network architectures for both KAN and ChebPIKAN to assess how the number of hidden layers affects the performance. Throughout this study, we maintained consistent hyperparameters and training strategies for both networks to ensure a fair comparison. Each neural network maintains an identical architecture, with the same number of neurons per hidden layer.
@@
$\mathbf{u}_{\text{res}}=\left(\frac{\sqrt{(\mathbf{u}_{\text{pred}}-\mathbf{u}_{\text{true}})^{2}}}{\sqrt{||\mathbf{u}_{\text{true}}||}}\right)\times 100,$ (16)
where $\mathbf{u}_{\text{res}}$, $\mathbf{u}_{\text{true}}$, and $\mathbf{u}_{\text{pred}}$ denote the residual, ground truth, and prediction of solution variables, respectively; |||||$\cdot$||| | ⋅ | | is the overall Euclidean-norm value.
@@ III.1 One-dimensional Allen–Cahn Equation
Allen–Cahn (AC) equation is a pivotal partial differential equation and widely applied in materials science, physical chemistry, and geometry. It effectively models the evolution of interfaces during phase transitions with a particular focus on single-phase transitions. This study uses the AC equation to assess the network’s capability in handling one-dimensional nonlinear problems with the reference solution derived from CFD simulations.
@@
For this study, we consider a computational domain of $[-1,1]$ and a time interval of $[0,1]$. In this domain, 225 ground truth data are randomly sampled for the training set. For ChebPIKAN, 800 PDE points are randomly selected to calculate the physical loss. Adam optimizer is used due to its efficiency; $10^{4}$ epoch iterations, an initial learning rate of 0.01, and a correction coefficient of 0.95 were used. The learning rate is adjusted downward every 10 steps once the loss curve stabilized.
It is noteworthy that, for the sake of descriptive convenience during the analysis, different numbers of hidden layers are used to represent distinct neural networks. However, the number of hidden layers, as a parameter, does not fundamentally affect the performance of neural networks. In essence, the underlying distinction lies in the variation of the number of trainable parameters. Within the same network architecture, the fitting capacity of the network is positively correlated with the number of trainable parameters. Theoretically, given sufficient training, a greater number of trainable parameters enables the neural network to model more complex problems. In addition, the conventional fully-connected Artificial Neural Networks (ANN) and PINN also serve as comparative neural network models and are trained accordingly.
@@
Figures 2 reveals that all KANs exhibit overfitting, which is particularly pronounced in those with three hidden layers. This finding suggests that KANs have suboptimal extrapolation performance. By contrast, ChebPIKANs, which incorporate physical information, effectively mitigate overfitting and notably reduce the testing loss associated with four hidden layers. By integrating physical information, ChebPIKANs enhance the networks’ ability to capture the essential characteristics of the model and consequently improve the extrapolation performance. In addition, ChebPIKANs consistently show lower testing loss values than KANs, which aligns with our expectation that the inclusion of physical information strengthens the network performance. We also observe that the training loss and PDE loss remain stable, which confirms that this integration helps prevent overfitting. Ultimately, the most effective ChebPIKAN configuration, which features four hidden layers, achieves loss convergence at approximately $10^{-3}$. Notably, when implementing fully-connected architectures with comparable parameter counts, both conventional ANNs and baseline PINNs exhibit significantly larger training loss values than the two proposed architectures. For PINNs, this training loss metric specifically quantifies only the data-matching component, explicitly excluding any physical constraint terms. The observed performance gap demonstrates these traditional architectures’ intrinsic representational limitations when constrained to simple topological designs.
@@
Figure 3 presents the error contours demonstrating that ChebPIKANs incorporate physical information to achieve more accurate predictions than KANs, as evidenced by the significantly reduced error distribution and magnitude. While KANs exhibit deteriorating prediction performance with increasing hidden layers due to heightened parameter complexity, ChebPIKANs maintain stability and shows steady accuracy improvement, reaching optimal performance with four hidden layers. In contrast, both ANNs and PINNs struggle to extract comprehensive flow field information from real data. Notably, the single-hidden-layer PINN fails to discern latent patterns in sparse data and becomes entirely guided by physics-informed constraints. Although a zero-velocity field technically satisfies the investigated AC equation, this trivial universal solution remains physically meaningless.
@@
For clarity, “$\textup{N}_{\textup{h}}$” in subsequent figures and tables denotes the number of hidden layers.
@@
Table 2 summarizes the average residuals for KANs, ChebPIKANs, ANNs and PINNs across different hidden layer configurations. ChebPIKANs consistently demonstrate lower mean residuals than KANs; in particular, with four hidden layers, KAN shows increased residuals, but ChebPIKAN does not. As comparative models, both ANNs and PINNs show suboptimal overall performance. However, in terms of the number of trainable parameters, the neural network enhanced by Chebyshev polynomials, as shown in Eq. (7), has only one trainable parameter $\omega$ between every two neuron nodes in adjacent network layers. Despite this, it learns significantly more effective information, indicating that the new algorithm effectively improves the information storage capacity per parameter. This is of great significance for large-scale models, as it can enhance information storage efficiency or reduce the number of network parameters, thereby facilitating the transfer of neural networks. In the study of the Allen–Cahn equation, sparse real data are used. Both ANNs and PINNs are entirely unable to effectively predict the flow field with a small-parameter network structure. In particular, the PINNs perform even worse due to the insufficient descriptive capacity of the small-parameter network, failing to effectively learn the physical equations.
We explored the effects of various neural networks structures on the network performance by configuring four different network architectures, each of which contained 1-4 hidden layers. Standard ANNs and PINNs with relatively simple architectures exhibit limited applicability to the problem under investigation. Theoretically, deeper networks have stronger nonlinear representative capabilities. However, with four hidden layers, KAN faces challenges due to an increase in parameters, which lead to poor training performance. Conversely, ChebPIKANs exhibit a steady decrease in loss with more hidden layers, which indicate that the integration of physical information enhance the networks’ learning capability. The improved overall nonlinear processing ability demonstrates the benefits of incorporating hidden layers into the architecture.
@@ III.2 One-dimensional Burgers Equation
$Los{s_{PDE}}=\nu\frac{\partial^{2}u}{\partial x^{2}}-\frac{\partial u}{\partial t}-u\frac{\partial u}{\partial x}.$ (22)
For the neural network, the inputs of Burgers equation are position $x$ and time $t$, and the output is fluid velocity $u$. The domain and random sampling strategies for the Burgers equation align with those used for the Allen–Cahn equation, and we consistently applied the Adam optimization strategy.
@@
The Optimal loss values in Figures 6 reveal that the final loss is less than ideal: most network configurations achieve a loss of approximately $10^{-2}$, close to $10^{-1}$. KANs’ overfitting is more pronounced in this context than in the Allen–Cahn equation. ChebPIKANs show resilience to overfitting, except with four hidden layers. Burgers equation presents substantial challenges due to its strong nonlinear phenomena and potential for shock wave generation at low viscosity coefficients. Consequently, when the network architecture deepens, the prediction accuracy significantly improves. ChebPIKANs show clear advantages with three hidden layers because this depth enhances its nonlinear processing capabilities; however, the structure with four hidden layers suffers from the complications of having more learning parameters. The training loss and PDE loss curves exhibit a high degree of alignment, which indicates that the inclusion of physical information mitigates overfitting in ChebPIKANs and consequently enhances the generalization. Thus, the network effectively learns the characteristics of the partial differential equations involved. Similar to studies on the Allen-Cahn equation, ANNs and PINNs with simple network architectures remain insufficient for addressing this class of problems.
@@ III.3 Two-dimensional Helmholtz Equation
In physical applications such as heat conduction and wave equations, $-u_{xx}$ and $u_{yy}$ are the changes in local curvature or temperature in $x$ and $y$ directions, respectively, which reflects the system’s diffusion or propagation. Meanwhile, $k_{0}^{2}u$ typically affects the system’s equilibrium state and acts as a restoring force or potential energy influence, whereas $f$ is an externally driven source term that represents the external influence or force applied in space and can be understood as the distribution of external stimuli or sources.
@@ III.4 Two-dimensional Kovasznay Flow
where $\lambda=\frac{1}{2\nu}-\sqrt{\frac{1}{4\nu^{2}}+4\pi^{2}}$.
This equation is a simplified form of Navier–Stokes equation in fluid dynamics and describes the conservation of momentum in fluid motion. Here, $Re$ is the Reynolds number, which characterizes the relative importance of inertia and viscous forces in flow. This item reflects the momentum diffusion and dissipation caused by shear in the fluids.
In our analysis, we define the spatial domain as $x\in[0,1]$ and $y\in[0,1]$. The same random sampling and optimization strategies are used as previously mentioned. The neural network model for Kovasznay flow uses two inputs ($x$ and $y$) and three outputs: pressure $p$, velocity $u$, and velocity $v$. In the ChebPIKANs frameworks, the physical losses associated with Kovasznay flow are expressed as
@@
For clarity in illustrating the performance of the four neural networks, Figure 15 presents a comparison of error contours exclusively for networks with three hidden layers. The results demonstrate that ANNs and PINNs with smaller architectures remain ineffective in handling physical field problems with sparse real-world data. In contrast, both KANs and ChebPIKANs exhibit strong predictive performance for the u-vector. ChebPIKANs, which are enhanced by additional physical information, exhibit virtually no prediction error. By contrast, KANs show a slight error at the boundary of the strong-shear region on the left. Since the magnitude of $v$ is smaller than that of the dominant $u$ and its solution exhibits a more complex form compared to both $u$ and $p$, both KANs and ChebPIKANs demonstrate lower predictive accuracy for $v$ than for $u$ and $p$. However, after incorporating boundary constraints, ChebPIKANs show a significant reduction in boundary errors, resulting in markedly lower overall errors than KANs. The predictive performance for $p$ is comparable between the two methods.
@@ III.5 Two-dimensional Unsteady Navier–Stokes Equation
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}=-\frac{\partial p}{\partial x}+\nu\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right),$ (33)
$\frac{\partial v}{\partial t}+u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}=-\frac{\partial p}{\partial y}+\nu\left(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\right).$ (34)
@@
$\begin{split}Los{s_{PDE}}=\,&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}+\frac{\partial p}{\partial x}-\nu\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right)\\ +\,&\frac{\partial v}{\partial t}+u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}+\frac{\partial p}{\partial y}-\nu\left(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\right)\\ +&\frac{\partial u}{\partial x}+\frac{\partial v}{\partial y}.\end{split}$ (36)

### v3

@@ II.3 Physics-informed KAN Network
For each physical equation, we design a varying number of hidden layers from shallow to deep to compare and select the optimal number of hidden layers. The Adam optimizer is used for the parameter adjustment. Specific parameters are presented in Table 1, where ILR is the initial learning rate, and $Data_{train}$ and $Data_{test}$ are the number of samples that have been used for training and testing, respectively. It should be noted that, to ensure fairness in neural network performance comparisons and to eliminate potential biases introduced by stochastic factors during training, we deliberately disabled the automatic pruning function. This guarantees that each neural network maintains a predictable number of trainable parameters throughout the experiments.
@@
[…]
@@
The two-dimensional input flow conditions are randomly sampled from a database of ground truth at a resolution of $\times15$ to form the training dataset, for more complex 2D cases like cavity flow, $\times25$ ground truth data is utilized as the training dataset. While such low-resolution data theoretically cannot fully describe the characteristics of the flow field, it more closely aligns with the sparsely discrete data obtained from real experiments through sensors. This presents a significant challenge to the inferential capabilities of the neural network, it forces the neural network to develop robust feature extraction capabilities under information-deficient conditions, better preparing it for practical deployment scenarios. For three-dimensional flow conditions, the amount of real data used for network training is appropriately increased.
@@ III Physical Models and Training Process of ChebPIKAN Models
For each equation, we explored six different network architectures for both KAN and ChebPIKAN to assess how the number of hidden layers affects the performance. Throughout this study, we maintained consistent hyperparameters and training strategies for both networks to ensure a fair comparison. Each neural network maintains an identical architecture, with the same number of neurons per hidden layer and polynomials of fixed 10th order.
@@
[…]
where $\mathbf{u}_{\text{res}}$, $\mathbf{u}_{\text{true}}$, and $\mathbf{u}_{\text{pred}}$ denote the residual, ground truth, and prediction of solution variables, respectively; $||\cdot||$ is the overall Euclidean-norm value.
@@ III.1 One-dimensional Allen–Cahn Equation
Allen–Cahn (AC) equation is a pivotal partial differential equation and is widely applied in materials science, physical chemistry, and geometry. It effectively models the evolution of interfaces during phase transitions with a particular focus on single-phase transitions. This study uses the AC equation to assess the network’s capability in handling one-dimensional nonlinear problems with the reference solution derived from CFD simulations.
@@
For this study, we consider a computational domain of [1,1][-1,1] and a time interval of $[0,1]$. In this domain, 225 ground truth data are randomly sampled for the training set. For ChebPIKAN, 800 PDE points are randomly selected to calculate the physical loss. Adam optimizer is used due to its efficiency; $10^{4}$ epoch iterations, an initial learning rate of 0.01, and a correction coefficient of 0.95 were used. The learning rate is adjusted downward every 10 steps once the loss curve stabilized.
It is noteworthy that, for the sake of descriptive convenience during the analysis, different numbers of hidden layers are used to represent distinct neural networks. However, the number of hidden layers, as a parameter, does not fundamentally affect the performance of neural networks. In essence, the underlying distinction lies in the variation of the number of trainable parameters. Within the same network architecture, the fitting capacity of the network is positively correlated with the number of trainable parameters. Theoretically, given sufficient training, a greater number of trainable parameters enables the neural network to model more complex problems. In addition, the conventional fully-connected artificial neural networks (ANN) and PINN also serve as comparative neural network models and are trained accordingly. Notably, both PINN and ChebPIKAN implementations adopt identical physical information loss terms and consistently maintain the same weighting coefficients $\lambda$.
@@
Figure 2 reveals that all KANs exhibit overfitting, which is particularly pronounced in those with three hidden layers. This finding suggests that KANs have suboptimal extrapolation performance. By contrast, ChebPIKANs, which incorporate physical information, effectively mitigate overfitting and notably reduce the testing loss associated with four hidden layers. By integrating physical information, ChebPIKANs enhance the networks’ ability to capture the essential characteristics of the model and consequently improve the extrapolation performance. In addition, ChebPIKANs consistently show lower testing loss values than KANs, which aligns with our expectation that the inclusion of physical information strengthens the network performance. We also observe that the training loss and PDE loss remain stable, which confirms that this integration helps prevent overfitting. Ultimately, the most effective ChebPIKAN configuration, which features four hidden layers, achieves loss convergence at approximately $10^{-3}$. Notably, when implementing fully-connected architectures with comparable parameter counts, both conventional ANNs and baseline PINNs exhibit significantly larger training loss values than the two proposed architectures. For PINNs, this training loss metric specifically quantifies only the data-matching component, explicitly excluding any physical constraint terms. The observed performance gap demonstrates these traditional architectures’ intrinsic representational limitations when constrained to simple topological designs.
@@
Figure 3 presents the error contours demonstrating that ChebPIKANs incorporate physical information to achieve more accurate predictions than KANs, as evidenced by the significantly reduced error distribution and magnitude. While KANs exhibit deteriorating prediction performance with increasing hidden layers due to heightened parameter complexity, ChebPIKANs maintain stability and show steady accuracy improvement, reaching optimal performance with four hidden layers. In contrast, both ANNs and PINNs struggle to extract comprehensive flow field information from real data. Notably, the single-hidden-layer PINN fails to discern latent patterns in sparse data and becomes entirely guided by physics-informed constraints. Although a zero-velocity field technically satisfies the investigated AC equation, this trivial universal solution remains physically meaningless.
@@
For clarity, $\textup{N}_{\textup{h}}$” in subsequent figures and tables denotes the number of hidden layers.
@@
Table 2 summarizes the average residuals for KANs, ChebPIKANs, ANNs and PINNs across different hidden layer configurations. ChebPIKANs consistently demonstrate lower mean residuals than KANs, in particular, with four hidden layers, KAN shows increased residuals, but ChebPIKAN does not. As comparative models, both ANNs and PINNs show suboptimal overall performance. However, in terms of the number of trainable parameters, the neural network enhanced by Chebyshev polynomials, as shown in Eq. (7), has only one trainable parameter $\omega$ between every two neuron nodes in adjacent network layers. Despite this, it learns significantly more effective information, indicating that the new algorithm effectively improves the information storage capacity per parameter. This is of great significance for large-scale models, as it can enhance information storage efficiency or reduce the number of network parameters, thereby facilitating the transfer of neural networks. In the study of the Allen–Cahn equation, sparse real data are used. Both ANNs and PINNs are entirely unable to effectively predict the flow field with a small-parameter network structure. In particular, the PINNs perform even worse due to the insufficient descriptive capacity of the small-parameter network, failing to effectively learn the physical equations.
We explored the effects of various neural network structures on the network performance by configuring four different network architectures, each of which contained 1-4 hidden layers. Standard ANNs and PINNs with relatively simple architectures exhibit limited applicability to the problem under investigation. Theoretically, deeper networks have stronger nonlinear representative capabilities. However, with four hidden layers, KAN faces challenges due to an increase in parameters, which lead to poor training performance. Conversely, ChebPIKANs exhibit a steady decrease in loss with more hidden layers, which indicate that the integration of physical information enhance the networks’ learning capability. The improved overall nonlinear processing ability demonstrates the benefits of incorporating hidden layers into the architecture.
@@ III.2 One-dimensional Burgers Equation
$Los{s_{PDE}}=\frac{0.01}{\pi}\frac{\partial^{2}u}{\partial x^{2}}-\frac{\partial u}{\partial t}-u\frac{\partial u}{\partial x}.$ (22)
For the neural network, the inputs of Burgers equation are position $x$ and time $t$, and the output is fluid velocity $u$. The domain and random sampling strategies for the Burgers equation align with those used for the Allen–Cahn equation, and we consistently apply the Adam optimization strategy.
@@
The optimal loss values in Figure 6 reveal that the final loss is less than ideal, most network configurations achieve a loss of approximately $10^{-2}$, close to $10^{-1}$. KANs’ overfitting is more pronounced in this context than in the Allen–Cahn equation. ChebPIKANs show resilience to overfitting, except with four hidden layers. Burgers equation presents substantial challenges due to its strong nonlinear phenomena and potential for shock wave generation at low viscosity coefficients. Consequently, when the network architecture deepens, the prediction accuracy significantly improves. ChebPIKANs show clear advantages with three hidden layers because this depth enhances its nonlinear processing capabilities; however, the structure with four hidden layers suffers from the complications of having more learning parameters. The training loss and PDE loss curves exhibit a high degree of alignment, which indicates that the inclusion of physical information mitigates overfitting in ChebPIKANs and consequently enhances the generalization. Thus, the network effectively learns the characteristics of the partial differential equations involved. Similar to studies on the Allen-Cahn equation, ANNs and PINNs with simple network architectures remain insufficient for addressing this class of problems.
@@ III.3 Two-dimensional Helmholtz Equation
In physical applications such as heat conduction and wave equations, $u_{xx}$ and $u_{yy}$ are the changes in local curvature or temperature in $x$ and $y$ directions, respectively, which reflects the system’s diffusion or propagation. Meanwhile, $k_{0}^{2}u$ typically affects the system’s equilibrium state and acts as a restoring force or potential energy influence, whereas $f$ is an externally driven source term that represents the external influence or force applied in space and can be understood as the distribution of external stimuli or sources.
@@ III.4 Two-dimensional Kovasznay Flow
where $\lambda=\frac{Re}{2}-\sqrt{\frac{Re^{2}}{4}+4\pi^{2}}$.
The Kovasznay flow is a specific form of the NS equation and thus inherently satisfies the constraints of the NS equation. In our analysis, we define the spatial domain as $x\in[0,1]$ and $y\in[0,1]$. The same random sampling and optimization strategies are used as previously mentioned. The neural network model for Kovasznay flow uses two inputs ($x$ and $y$) and three outputs: pressure $p$, velocity $u$, and velocity $v$. In the ChebPIKANs frameworks, the physical losses associated with Kovasznay flow are expressed as
@@
For clarity in illustrating the performance of the four neural networks, Figure 15 presents a comparison of error contours exclusively for networks with three hidden layers. The results demonstrate that ANNs and PINNs with smaller architectures remain ineffective in handling physical field problems with sparse real-world data. In contrast, both KANs and ChebPIKANs exhibit strong predictive performance for the u-vector. ChebPIKANs, which are enhanced by additional physical information, exhibit virtually no prediction error. By contrast, KANs show a slight error at the boundary of the strong-shear region on the left. Since the magnitude of $v$ is smaller than that of $u$ and the solution of $v$ exhibits a more complex form, both KANs and ChebPIKANs show lower predictive accuracy for $v$. However, after incorporating boundary constraints, ChebPIKANs show a significant reduction in boundary errors, resulting in markedly lower overall errors than KANs. The predictive performance for $p$ is comparable between the two methods.
@@ III.5 Two-dimensional Unsteady Navier–Stokes Equation
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}=-\frac{\partial p}{\partial x}+\frac{1}{Re}\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right),$ (33)
$\frac{\partial v}{\partial t}+u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}=-\frac{\partial p}{\partial y}+\frac{1}{Re}\left(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\right),$ (34)
@@
$\begin{split}Los{s_{PDE}}=\,&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}+\frac{\partial p}{\partial x}-\frac{1}{Re}\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right)\\ +\,&\frac{\partial v}{\partial t}+u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}+\frac{\partial p}{\partial y}-\frac{1}{Re}\left(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\right)\\ +&\frac{\partial u}{\partial x}+\frac{\partial v}{\partial y}.\end{split}$ (36)